If David’s brother is 3 years older, and the sum of their ages is 17, how old is David?
Answer by indra89811(24)

Let David's age =x
David's brother's age=x+3
By question, x + x + 3 = 17
2x+3=17
2x=17-3
2x=14
x=7
Hence, David is 7 years old.

-------------------Peace out, hun :)-------------------------